Determine the functions f: R -> R are onto, f(x) = |x| + x



1
Expert's answer
2022-06-13T17:51:31-0400
"x<0: f(x)=-x+x=0"

"x\ge0: f(x)=x+x=2x"

Then


"f(x)= \begin{cases}
 0& x<0 \\
 2x & x\ge0
\end{cases}"

Domain: "(-\infin, \infin)"

Range: "[0, \infin)"


The function "f: R\to R" is not onto because there is no "x\in \R" with "|x|+x=-1," for instance.